OpenStudy (anonymous):

Find particular solution to differential equation. 5y'(t) + 10y(t) = 2e^-t

OpenStudy (anonymous):

Are you familiar with method of integrating factor?

OpenStudy (anonymous):

I used to be (; I will look it up.

OpenStudy (anonymous):

I'll get you started! \[p=10\]\[P= \int\limits 10~dt\]\[\mu = e^P\]

OpenStudy (anonymous):

Oops, I made a stupid mistake. Coefficient in front of the second order derivative must be 1! \[\large p=5\]\[\large P=\int\limits 5~dt\]\[\large\mu=e^{5t}\]\[\large e^{5t}(y'+2y)=\frac{ 2 }{ 5 }e^{-t}e^{5t}\]\[\large (e^{5t}y)'=\frac{ 2 }{ 5 }e^{4t}\]

OpenStudy (anonymous):

Continue to solve for y!

OpenStudy (anonymous):

Shouldn't it be e^2t * y = Integral (e^2t)(2/5)(e^-t) so y = (2/5)e^-t

OpenStudy (anonymous):

Yes, I'm sorry again, I can't even simple math XD I was up 'till 6am studying for an engineering statics midterm I had today. But you are correct! \(\large (e^{2t}y)'=\frac{ 2 }{ 5 }e^{t}\) Don't forget the \(+C\)! :)

OpenStudy (anonymous):

Thanks for your time!!

OpenStudy (anonymous):

You are most certainly welcome! X)